OpenStudy (anonymous):

dy/dx = x+y/y-x' , y(1)=2 @zepdrix

OpenStudy (turingtest):

x'=dx/dy ??

OpenStudy (anonymous):

i got a couple steps in and lost it

OpenStudy (turingtest):

I am asking you, what does x' represents here? or is that a typo?

OpenStudy (anonymous):

it represents x prime

zepdrix (zepdrix):

That doesn't make sense patty :c

zepdrix (zepdrix):

The prime would suggest it's a `derivative of x`. Derivative with respect to what other variable? With respect to y? So it does represent dx/dy?

zepdrix (zepdrix):

Or that's all the info you were given? c:

OpenStudy (anonymous):

screw x prime and say the ' was a typo

zepdrix (zepdrix):

lol XD

OpenStudy (anonymous):

z anything?

OpenStudy (turingtest):

let v=y-x then v'=y'-1 -> y'=v'+1 we then have v'+1=(v+2x)/v try solving that, then subbing back in for v in terms of y after integration

OpenStudy (anonymous):

i think i did something wrong

OpenStudy (turingtest):

can you try to show some of what you did?